# Safe Control using Occupancy Grid Map-based Control Barrier Function (OGM-CBF)

### v2

@@ III-A Occupancy Grid Map-based Control Barrier Function (OGM-CBF)
$h(\boldsymbol{x})$ $=\Phi_{s}(\boldsymbol{x},m)+l_{s}+l_{a}cos(\eta)$ (7)
@@
  • $\Phi_{s}(\boldsymbol{x},m)=T(\phi(\boldsymbol{x},M_{b}(m)))$, and $T:\mathbb{R}\rightarrow\mathbb{R}$ is a scalar odd function for shaping and smoothing. In other words, $\Phi_{s}$ is a smoothened signed distance function of the set $\mathbb{X}_{s}$, evaluated at the robot’s body frame position $p$ in the map, noting that $p$ is part of the state vector $\boldsymbol{x}$.
  • $\eta$ is the angle between vectors $\hat{\boldsymbol{x}}$, and $\nabla_{p}\Phi_{s}(\boldsymbol{x},m)$, where $\hat{\boldsymbol{x}}$ is the unit vector of the first coordinate of the body frame, $\nabla_{p}\Phi_{s}(\boldsymbol{x},m)$ is gradient of $\Phi_{s}(\boldsymbol{x},m)$.
@@
Note that in above equations, all the quantities such as body frame and map frame, and accordingly $\Phi_{s}$ are defined in the same coordinate frame, for instance, the robot ego-centric frame.
Figure 1: Top-view snapshot of the environment in CARLA. The second row shows the constructed occupancy grid map, and the last row shows $\Phi_{s}(\boldsymbol{x},m)$ and its gradient field in the map frame for all the states in the map.
Next, we provide some intuition of the proposed CBF in (7). The proposed CBF function comprises two main components: i) the effect of the robot’s distance to obstacles, and ii) the angle between the robot’s heading and the gradient of the SDF function $\Phi_{s}$.
In the first component, $\Phi_{s}(\boldsymbol{x},m)$, map $m$ captures the latest location, shape, and size of the obstacles within the map at each time step, as outlined in (4). In the map, each occupied cell indicates an obstacle. Although effective, this map lacks smooth transitional values. In Fig. 1, the top row displays a top view of the scenario and obstacle status, while the second row depicts the resulting binary occupancy grid map. Function $\phi$ is smooth almost everywhere except at the curves where the closest obstacle region switches. Applying $T(.)$ to $\phi$ ensures smooth $\Phi_{s}$ everywhere. The last row in Fig. 1 illustrates the smooth $\Phi_{s}$ in the map coordinate.
The second component, $l_{s}+l_{a}cos(\eta)$, involves calculating the cosine similarity between two vectors: $\hat{x}$, representing the robot heading, and $\nabla_{p}\Phi_{s}(\boldsymbol{x},m)$. Notice that $\nabla_{p}\Phi_{s}$ is the direction that $\Phi_{s}$ increases the most, thus moving in that direction will move the robot further away from obstacles. The gradient field depicted in the last row of Fig. 1 visualizes $\nabla_{p}\Phi_{s}(\boldsymbol{x},m)$ within the map frame. Notice that $l_{s}$ is a negative number and the value $-l_{s}-l_{a}\geq 0$ is the safety margin, a design choice.
@@
Set $S=\left\{\boldsymbol{x}\in\mathbb{X}\;|\;h(\boldsymbol{x})\geq 0\right\}$ with $h(\boldsymbol{x})$ defined in (7) is a forward-invariant set and $h(\boldsymbol{x})$ is a control barrier function for system (1), with $\boldsymbol{u}=[\mathrm{v},\boldsymbol{\omega}]^{T}$ as the control input.
@@
We will first show that $S\subseteq\mathbb{X}_{s}$. The condition $0<l_{a}\leq-l_{s}$ ensures that $\forall\boldsymbol{x}$ we have $l_{s}+l_{a}cos(\eta)\leq 0$. For any $\boldsymbol{x}\in S=\{\boldsymbol{x}:h(\boldsymbol{x})\geq 0\}$, we have $h(\boldsymbol{x})\geq 0$, therefore $\Phi_{s}(\boldsymbol{x})\geq 0$, and consequently $\phi(\boldsymbol{x})\geq 0$, since $T(.)$ is an odd function. Thus $\boldsymbol{x}\in\mathbb{X}_{s}=\{\boldsymbol{x}:\phi(\boldsymbol{x})\geq 0\}$, and we have shown the first part.
We will now show that there are control signals $\boldsymbol{u}$ such that (3) is satisfied. It is easy to see that $\dot{h}=\frac{\partial\Phi_{s}}{\partial x}\dot{p}+l_{a}\sin(\eta)\dot{\eta}$, which can be written as $\dot{h}=a(x)v+b(x)\boldsymbol{\omega}$, noting that $\dot{p}$ and $\dot{\eta}$ provide the $v$ term, while $\dot{\eta}$ provides the $\boldsymbol{\omega}$ term, assuming a kinematic model. Since $(v,\boldsymbol{\omega})$ can be made zero, condition (3) can be satisfied and thus $h(\boldsymbol{x})$ is a control barrier function for (1), and $S$ is an invariant set. ∎
@@
Considering the control system in (1), with control inputs $\boldsymbol{u}=[\mathrm{v},\boldsymbol{\omega}]^{T}$, the control policy outlined in (8) ensures safety across all trajectories originating within $S$.
$\pi_{ogm-cbf}=\{\boldsymbol{u}\in\mathbb{U}\;|\;\dot{h}(\boldsymbol{x},\boldsymbol{u})\geq-\alpha(h(\boldsymbol{x}))\}$ (8)
@@
It is important to recognize that the efficacy of the safety controller is influenced by the configuration of the function $T(.)$, $\alpha(.)$, and the values of $l_{s}$, and $l_{a}$. which can be adjusted according to the model of the system, the dimensions of the map, and the environment.
Note that since the union of all occupied cells is considered in the function $h(\boldsymbol{x})$, and the distance to the closest occupied cell is being calculated in this function, the OGM-CBF automatically will find the arbitrary safe set $S$, that considers all the detected obstacles. This means, in contrast to other methods, one CBF construction is enough at each timestep to guarantee staying in the safe set $S$, which makes this method computationally efficient.
@@
Figure 2: Top-view snapshot of simulation in the CARLA environment. The blue car is a self-driving vehicle using OGM-CBF to navigate safely. The orange dashed line is the path of the autonomous car. The second row shows the upper half of the ego-centric local occupancy grid map denoted as $m$, and the last line illustrates $\Phi_{s}(\boldsymbol{x},m)$ and its gradient field in the map frame for all the states in the map.
III-B Including Robot shape in OGM-CBF
OGM-CBF is capable of considering the robot’s size and shape using established inflation methods on occupancy grid maps. For example, the robot can be modeled as a union of multiple spheres, similar to that described in [26]. In addition, the occupied cells in the map can be inflated by the radius of the spheres. Then the designed CBF can be analyzed at the centers of these spheres.
One can also choose number of points at the boundary of the robot to be avoided, with one constraint per point in OGM-CBF. Having enough points, the OGM-CBF ensures that the entire boundary of the robot remains within the safe set $S$.
III-C Control Lyapunov Function (CLF)
@@
$\inf_{\boldsymbol{u}\in\mathbb{U}}[\dot{V}(\boldsymbol{x},\boldsymbol{u})=\frac{\partial V}{\partial\boldsymbol{x}}(f(x)+g(x)\boldsymbol{u})]\leq-\gamma(V(\boldsymbol{x}))$ (9)
@@
III-D Integrating CLF and OGM-CBF
In order to fulfill the criteria for both safety and stability, the proposed control policy should represent the intersection of two preceding policies, denoted as $\pi^{*}=\pi_{ogm-cbf}\cap\pi_{clf}$. In instances where this intersection is empty, prioritization of safety becomes imperative. Consequently, the following constrained optimization problem can be defined to find an optimal control input that simultaneously satisfies the constraints related to safety and stability, respectively as hard and soft constraints.
$\boldsymbol{u}^{*}$ $=\underset{\boldsymbol{u}}{\mathrm{argmin}}\;J(\boldsymbol{u})$ (11a)
[…]
@@ III-D Integrating CLF and OGM-CBF
$\boldsymbol{u}_{lb}$ $\leq\boldsymbol{u}\leq\boldsymbol{u}_{ub}$ (11d)
where $J$ represents the cost function, while $\boldsymbol{u}_{lb}$ and $\boldsymbol{u}_{ub}$ are the lower and upper bounds of the control input, respectively. Additionally, $\delta$ is identified as the relaxation parameter, which is engaged when the constraints are in conflict, leading to an absence of elements in the intersection $\pi_{ogm-cbf}\cap\pi_{clf}$. Under such circumstances, the hard constraint (related to safety) is maintained, whereas the soft constraint (related to stability) is moderated. It is customary to set a quadratic cost $J(\boldsymbol{u})$, which renders the optimization a Quadratic Programs (QP), since the constraints are linear in control $u$.
In the Algorithm (1) steps of deployment of the proposed map-based safe control is expressed.
Input: sensor data: $z$, state of the system: $\boldsymbol{x}$
@@
2
3      receive sensor data $z$ and state of the system $\boldsymbol{x}$
4       generate occupancy grid map $m$
5
6      generate $\Phi_{s}(\boldsymbol{x},m)$ and its gradient
7       calculate $h(\boldsymbol{x})$, and $\dot{h}(\boldsymbol{x},\boldsymbol{u})$ for the current state of the robot
@@
IV function shaping and smoothing
In this section, we aim to explore the properties of the function $T(.)$ used in generating $h(\boldsymbol{x})$ as a control barrier function. Given that the signed distance function is continuously differentiable almost everywhere, as discussed in section II-C, the design of $T(\phi(\boldsymbol{x},M_{b}(m)))$ ensures that $h(\boldsymbol{x})$ maintains continuous differentiability across the entire state space. Notably, $\phi(.)$ may lack differentiability at non-smooth points or centers of the set $\Omega$. To address this, $T(\phi(\boldsymbol{x},M_{b}(m)))$ translates the grid map values of $\phi(\boldsymbol{x},M_{b}(m))$ to the state space and employs polynomial interpolation $p(.)$, thus guaranteeing continuous differentiability at all points in the state space. This approach ensures that $h(\boldsymbol{x})$ is a continuously differentiable function, which is required as specified in (2). As stated before, care needs to be taken as $T(.)$ needs to be an odd function.
Moreover, $T(.)$ enhances the control over the shape of $h(\boldsymbol{x})$ by adjusting the gradient behaviors and scaling the values of $\phi(\boldsymbol{x},M_{b}(m))$ through non-linear modifications prior to polynomial interpolation. This adjustment is tailored to fit the dynamics of the system described in (1).
@@
Figure 4: 2D snapshot of the robot during the experiment (left), the constructed occupancy grid map using onboard sensors (middle), and $\Phi_{s}(\boldsymbol{x},m)$ and its gradient field in the map frame for all the states in the map (right). The orange line shows the path of the robot and the red gradients are the taken gradient values to the algorithm during the experiment.

### v3

@@ III-A Occupancy Grid Map-based Control Barrier Function (OGM-CBF)
$h(\boldsymbol{x})=\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})+l_{s}+l_{a}\hat{\boldsymbol{x}}\cdot\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$ (8)
@@
  • $\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})=T(\phi(\boldsymbol{p},\mathbb{X}_{s}))$, and $T:\mathbb{R}^{j}\rightarrow\mathbb{R}$ is a odd function for shaping and smoothing, where $j$ is the number of points from the OGM used for interpolation. In other words, $\Phi_{s}$ is a smoothened signed distance function of the set $\mathbb{X}_{s}$, evaluated at the robot’s body frame position $\boldsymbol{p}$.
  • $\hat{\boldsymbol{x}}\cdot\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$ is the dot product of $\hat{\boldsymbol{x}}$, and $\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$, where $\hat{\boldsymbol{x}}$ is the robot’s heading unit vector, $\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$ is gradient of $\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$.
@@
  • Smoothing/Shaping function $T$ is defined such that $||\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})||\leq 1$. This ensures that $\forall\boldsymbol{x}\in X,l_{s}+l_{a}\hat{\boldsymbol{x}}\cdot\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})\leq 0,$
Note that in above equations, all the quantities such as $\boldsymbol{p}$, $\hat{\boldsymbol{x}}$, and accordingly $\Phi_{s}$ are defined in the same coordinate frame, for instance, the world frame, and $\boldsymbol{p}$, $\hat{\boldsymbol{x}}$ are derived from state of the system as described in (1).
Figure 1: The process of $\Phi_{s}$ construction. a) Top-view snapshot of the environment in CARLA b) the constructed occupancy grid map c) $\Phi_{s}$ and its gradient field in the map frame
Next, we provide some intuition of the proposed CBF in (8). The proposed CBF function comprises two main components: i) The first component captures the effect of the robot’s distance to obstacles, and ii) the remaining captures the alignment of the robot’s heading vector and the gradient of $\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$. The larger the components, the safer the robot.
Note that the occupancy grid map $m$ is calculated in discrete positions (centers of the cells). Consequently, due to computational reason and grid representation of $m$, $\phi(.,\mathbb{X}_{s})$ is initially calculated in discrete grid points of the map, and then smoothly interpolated to approximate the SDF at robot position $\boldsymbol{p}$ (space of real numbers). Function $T$ represents this interpolation (smoothing) and a non-linearity to shape the SDF function. Recall that $|\nabla\phi|=1$, thus applying $T$ allows adjusting the gradient and scaling of $\phi(\boldsymbol{p},X_{s})$ to a desired form. This non-linearity can be applied via functions such as $tanh$. Since the norm of the gradient of $tanh$ is less than $1$, all required conditions are satisfied. Moreover, applying the smooth interpolation operator as described ensures the differentiability of $h(\boldsymbol{x})$. Fig. 1 illustrates the steps involved in constructing $\Phi_{s}$. In subfigure (a), the CARLA environment is shown in a top-view, which is mapped using a binary OGM in subfigure (b). Finally, $\Phi_{s}$ is constructed as $T(\phi(.,\mathbb{X}_{s}))$, as visualized in subfigure (c).
The second component, $l_{s}+l_{a}\hat{\boldsymbol{x}}\cdot\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$, involves calculating the cosine similarity between two vectors: $\boldsymbol{\hat{x}}$, representing the robot heading, and $\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$. Notice that $\nabla_{p}\Phi_{s}$ is the direction that $\Phi_{s}$ increases the most, thus moving in that direction will move the robot further away from obstacles. The gradient field depicted in Fig. 1 (c) visualizes $\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$ within the map frame. Notice that $l_{s}$ is a negative number and the value $-l_{s}-l_{a}\geq 0$ is the safety margin, a design choice. The roles of these coefficients become more clear in the proof of Theorem 1. Moreover, notice that without the second component, derivative of $h(\boldsymbol{x})$ will be only a function of linear velocity, that is, to maintain the safety, the system can only stop. To enable steering, the solution is to use a higher order CBF [23] thus additional complexity. By adding the second term, the angular velocity appears in the first derivative of $h(\boldsymbol{x})$, and allows also steering around obstacles. A similar idea was used in [24].
@@
Set $S=\left\{\boldsymbol{x}\in\mathbb{X}\;|\;h(\boldsymbol{x})\geq 0\right\}$ with $h(\boldsymbol{x})$ defined in (8) is a forward-invariant set and $h(\boldsymbol{x})$ is a control barrier function for system (1), with $\boldsymbol{u}=[\boldsymbol{v},\boldsymbol{\omega}]^{T}$ as the control input.
@@
We will first show that $S\subseteq\mathbb{X}_{s}$. The condition on $T$ and $0<l_{a}\leq-l_{s}$ ensure that $\forall\boldsymbol{x}$ we have $c\equiv l_{s}+l_{a}\hat{\boldsymbol{x}}\cdot\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})\leq 0$. For any $\boldsymbol{x}\in S=\{\boldsymbol{x}:h(\boldsymbol{x})\geq 0\}$, we have $h(\boldsymbol{x})\geq 0$, therefore $\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})=h(\boldsymbol{x})-c\geq 0$, and consequently $\phi(\boldsymbol{p},\mathbb{X}_{s})\geq 0$, since $T(.)$ is an odd function. Thus $\boldsymbol{x}\in\mathbb{X}_{s}=\{\boldsymbol{x}:\phi(\boldsymbol{p},\mathbb{X}_{s})\geq 0\}$, proving the first part.
We will now show that there are control signals $\boldsymbol{u}$ such that (3) is satisfied. We will first derive $\dot{h}(\boldsymbol{x},\boldsymbol{u})$. Time derivative of $h$ is
$\dot{h}$ $=\frac{d}{dt}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})+\frac{d}{dt}\left(l_{a}\hat{\boldsymbol{x}}\cdot\nabla_{p}\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})\right),$ (9)
and assuming the environment is stationary ($\mathbb{X}_{s}$ is not changing), we have
$\dot{h}=\nabla_{p}\Phi_{s}\cdot\dot{\boldsymbol{p}}+l_{a}\frac{d\hat{\boldsymbol{x}}}{dt}\cdot\nabla_{p}\Phi_{s}+l_{a}\hat{\boldsymbol{x}}\cdot\frac{d}{dt}\nabla_{p}\Phi_{s}.$ (10)
Note that $\hat{\boldsymbol{x}}=R\hat{\boldsymbol{e}}$, where $R$ is the rotation matrix expressing the body in the world frame, and $\hat{\boldsymbol{e}}=[1,0,0]$ the x-axis of the world frame. Given that $\dot{\boldsymbol{p}}=R\boldsymbol{v}$ and $\frac{d\hat{\boldsymbol{x}}}{dt}=R[\boldsymbol{\omega}]\hat{\boldsymbol{e}}$, the equation (10) can be written as $\dot{h}=a(\boldsymbol{x})\boldsymbol{v}+b(\boldsymbol{x})\boldsymbol{\omega}$. Since, $(\boldsymbol{v},\boldsymbol{\omega})$ can be made zero, condition (3) can be satisfied and thus $h(\boldsymbol{x})$ is a control barrier function for (1), and $S$ is an invariant set. ∎
@@
Considering the control system in (1), with control inputs $\boldsymbol{u}=[\boldsymbol{v},\boldsymbol{\omega}]^{T}$, the control policy outlined in (11) ensures safety across all trajectories originating within $S$.
$\pi_{ogm-cbf}=\{{(\boldsymbol{v},\boldsymbol{\omega})}\in\mathbb{U}_{s}\;|\;a(\boldsymbol{x})v+b(\boldsymbol{x})\omega\geq-\alpha(h(\boldsymbol{x}))\}$ (11)
@@
It is important to recognize that the efficacy of the safety controller is influenced by the design of the function $T(.)$, $\alpha(.)$, and the values of $l_{s}$, and $l_{a}$, which can be adjusted according to the model of the system, the dimensions of the map, and the environment.
@@
Figure 2: Top-view snapshot of the simulation in the CARLA environment. The black car is a self-driving vehicle using OGM-CBF to navigate safely. The orange dashed line represents the path of the autonomous car. The second row shows the upper half of the egocentric local occupancy grid map denoted as $m$, and the last line illustrates $\Phi_{s}$ and its gradient field in the map frame during the experiment.
III-B Control Lyapunov Function (CLF)
@@
$\inf_{\boldsymbol{u}\in\mathbb{U}}[\dot{V}(\boldsymbol{x},\boldsymbol{u})=\frac{\partial V}{\partial\boldsymbol{x}}(f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u})]\leq-\gamma(V(\boldsymbol{x}))$ (12)
@@
III-C Integrating CLF and OGM-CBF
To ensure both safety and stability, the proposed control policy is defined as the intersection $\pi^{*}=\pi_{ogm-cbf}\cap\pi_{clf}$. If the intersection is empty, safety is prioritized. This leads to the following constrained optimization problem, balancing safety and stability as hard and soft constraints, respectively.
$\boldsymbol{u}^{*}=\underset{\boldsymbol{u}}{\mathrm{argmin}}\;J(\boldsymbol{u},\boldsymbol{x},\delta)$ (14a)
$\mathrm{s.t.:}\quad\dot{V}(\boldsymbol{x},\boldsymbol{u})\leq\gamma(V(\boldsymbol{x}))+\delta$ (14b)
@@ III-C Integrating CLF and OGM-CBF
$\boldsymbol{u}_{lb}\leq\boldsymbol{u}\leq\boldsymbol{u}_{ub},$ (14d)
where $J$ represents the cost function, while $\boldsymbol{u}_{lb}$ and $\boldsymbol{u}_{ub}$ are the lower and upper bounds of the control input, respectively. Additionally, $\delta$ is introduced as the relaxation parameter, activated when constraints conflict, resulting in an empty intersection between $\pi_{ogm-cbf}$ and $\pi_{clf}$. Under such circumstances, the hard constraint (related to safety) is maintained, whereas the soft constraint (related to stability) is moderated.
Since system (1) is affine in $u$, both $\dot{V}$ and $\dot{h}$ will become control-affine, more specifically
$\dot{V}(\boldsymbol{x},\boldsymbol{u})=a_{1}(\boldsymbol{x})+b_{1}(\boldsymbol{x})\boldsymbol{u}$ (15a)
$\dot{h}(\boldsymbol{x},\boldsymbol{u})=a_{2}(\boldsymbol{x})+b_{2}(\boldsymbol{x})\boldsymbol{u}$ (15b)
for some $a_{i}(\boldsymbol{x}),b_{i}(\boldsymbol{x}),i=1,2$. A quadratic cost function $J$ used in our experiments is
$J=(\boldsymbol{u}-\boldsymbol{u}_{ref}(\boldsymbol{x}))^{T}(\boldsymbol{u}-\boldsymbol{u}_{ref}(\boldsymbol{x}))+\delta^{2}$ (16a)
Since $J$ is quadratic and constraints are linear in $(\boldsymbol{u},\delta)$, the optimization (14) is a Quadratic Programs (QP). Algorithm (1) details the steps of deployment of the proposed OGM-CBF. In the algorithm, we assume that, given the observations, we can calculate state-dependent variables $\boldsymbol{u}_{ref}(\boldsymbol{x})$, $V(\boldsymbol{x})$, $h(\boldsymbol{x})$, $a_{i}(\boldsymbol{x}),b_{i}(\boldsymbol{x}),$ for $i=1,2$.
Input: observations $z$, $\boldsymbol{p}$, and $\hat{\boldsymbol{x}}$
@@
2       receive observations $z$, $\boldsymbol{p}$, and $\hat{\boldsymbol{x}}$
3       update occupancy grid map $m$ and safe set $\mathbb{X}_{s}$
4       calculate $\Phi_{s}(\boldsymbol{p};\mathbb{X}_{s})$ and its gradient
5       calculate $\boldsymbol{u}_{ref}(\boldsymbol{x})$, $V(\boldsymbol{x})$, $h(\boldsymbol{x})$, $a_{i}(\boldsymbol{x}),b_{i}(\boldsymbol{x}),i=1,2$
@@
Figure 4: Safe navigation of MiR100 using OGM-CBF. a) 2D snapshot of the robot during the experiment, b) the constructed occupancy grid map using onboard sensors, c) $\Phi_{s}$ and its gradient field in the map frame. The orange line shows the path of the robot and the red gradients represent the gradient values used by the algorithm during the experiment.
IV Qualitative Comparison
In this section, we clarify some of our contributions, compared to prior art.
Sensor Agnostic: OGM-CBF uses OGM to generate safe and unsafe sets. OGM can be constructed with sensors like LiDAR, radar, and cameras, making it adaptable to various sensors and environmental features. In contrast, [13, 11, 12] rely solely on one sensor modality, limiting their effectiveness in challenging conditions, such as poor lighting or reflective surfaces for RGB-D sensors and adverse weather conditions for LiDARs, where they are prone to extreme noise.
Spatial Awareness: OGM accumulates sensor data over time, providing a comprehensive and continuously updated environmental map for OGM-CBF, unlike V-CBF [13], which relies on instantaneous measurements. In [13], CBF’s awareness of the obstacles only relies on the current coverage of sensors (RGB-D cameras). In [11], previously constructed unsafe set is considered in CBF, while storing all the previous sensor measurements for online training of unsafe set detection, which is computationally expensive. To address this challenge, [12] leverages incremental online training with replay memory in SDF approximation.
Robustness: There are two sources of uncertainties: (i) perception sensors, and (ii) the robot pose/state estimation. OGM takes into account both uncertainties during the creation of the map. Moreover, when the localization is too uncertain, even if the map is created a priori with more accurate systems, to retain robust safety OGM-CBF provides the designer with possibility to tune $l_{a}$ and $l_{s}$ for an extra safety margin. Therefore, our method is robust in that sense, compared to [13] [12], which has no mechanism to handle sensor uncertainties. In [11] a robustness to sensor noise is provided in the SVM formulation.
Efficacy: In [13], [12], [25], [9], one constraint is needed per obstacle. In contrast, in OGM-CBF and similarly in [11], with the effective use of SDFs, a single constraint can account for all detected obstacles, making it more efficient for real-time deployment.